Encircle the odd one of the following 
 (a) (–9) × 5 × 6 × (–3) (b) 9 × (–5) × 6 × (–3) 
(c) (–9) × (–5) × (–6) × 3 (d) 9 × (–5) × (–6) × 3 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
(–9) × 5 × 6 × (–3) = 810
9 × (–5) × 6 × (–3) = 810
(–9) × (–5) × (–6) × 3 = -810
9 × (–5) × (–6) × 3 = 810
Hence, (–9) × (–5) × (–6) × 3 is the odd one.

Thanks